OpenStudy (love_333):

Help with two questions? Identify the vertex, axis of symmetry, maximum or minimum, and domain and range of each function. 6. F(x) = 10(x-1)^2 + 5 7. F(x) = 2(x+9)^2 - 4

OpenStudy (love_333):

If someone could please help me with these and also explain how to do them that'd be great.

OpenStudy (joshmorris):

hgiuk

OpenStudy (bonnieisflash1.0):

hello

OpenStudy (bonnieisflash1.0):

want to be friends @Love_333

OpenStudy (love_333):

Can you help me with this problem? @bonnieisflash1.0 @JoshMorris

OpenStudy (bonnieisflash1.0):

sure

OpenStudy (love_333):

I'm not sure how to do it or what its asking for

OpenStudy (bonnieisflash1.0):

what answer do you think it is

OpenStudy (love_333):

I have no idea, I don't really know what its asking me to do

OpenStudy (bonnieisflash1.0):

i can help if you want me to

OpenStudy (love_333):

That'd be great

OpenStudy (bonnieisflash1.0):

okay

OpenStudy (bonnieisflash1.0):

i think The parabola has equation y = -x^2 - 4x + 3. To find the vertex, complete the square. y - 3 = -x^2 - 4x y - 3 = -(x^2 + 4x) y - 3 - 4 = -(x^2 + 4x + 4) y = -(x + 2)^2 + 7 The vertex is (-2, 7). The axis of symmetry is x = -2. The maximum value is 7. The domain is {real numbers}. The range is {y: y ≤ 7}.

OpenStudy (bonnieisflash1.0):

i might be off track

OpenStudy (bonnieisflash1.0):

you can tell me if i'm off track

OpenStudy (love_333):

So how do I use that with my equations?

OpenStudy (bonnieisflash1.0):

\[f(x) = 10(x-1)^2 + 5\]

OpenStudy (bonnieisflash1.0):

this can probably help you

OpenStudy (love_333):

I stil don't really understand.. I think I figured out the domain and range though. Real numbers and \[f \ge5\]

OpenStudy (bonnieisflash1.0):

okay

OpenStudy (love_333):

Is that right? idk what im doing

OpenStudy (bonnieisflash1.0):

it could be right

OpenStudy (love_333):

vertex (1,5)?

OpenStudy (bonnieisflash1.0):

yep